OpenStudy (anonymous):

How do I derive this: 1/5^2

OpenStudy (anonymous):

\[\frac{1}{5^2}=\frac{1}{25}\] not sure what "derive" means in this context compute?

OpenStudy (loser66):

My answer : \((\dfrac{1}{5})^2\)